OpenStudy (kj4uts):

For a complex number written in polar form, it's representation is unique as long as the argument is restricted to [0, 2pi). True or False?

OpenStudy (anonymous):

False

OpenStudy (dumbcow):

counter-example: \[1+i = \sqrt{2}(\cos \frac{\pi}{4} +i \sin \frac{\pi}{4})\] \[1+i = -\sqrt{2}(\cos \frac{5\pi}{4} +i \sin \frac{5\pi}{4})\]

OpenStudy (anonymous):

see why this would be interesting to the mathematicians here!LOL